OpenStudy (anonymous):

(1-sec^2x)/(csc^2-1)=? worked solutions pleasee :)

OpenStudy (anonymous):

i believe its -1

OpenStudy (anonymous):

1-sec^2x= -tan^2x csc^2x-1= Cot^2x (-tan^2x)/(cot^2x)= -1

OpenStudy (anonymous):

thank you :)